rose and luke decided to take part in a twenty mile race. Luke runs at 5mph, while rose runs at 4mph. when Luke crosses the finish line, how far will rose be from finishing?
Answer by stanbon(75887)

rose and luke decided to take part in a twenty mile race.
Luke runs at 5mph,
Luke DATA:
rate = 5 mph ; distance = 20 miles ; time = 20/5 = 4 hrs
------------------------
while rose runs at 4mph.
Rose DATA:
rate = 4 mph ; distance = 20 miles; time = 20/4 = 5 hrs
----------------------------------------
when Luke crosses the finish line, how far will rose be from finishing?
---
Rose will be 1 hr. from the finish line.
Her distance from the finish line will be 4 mph * 1 hr = 4 miles.
======================
Cheers,
Stan H.
==============